OpenStudy (anonymous):

solve square root of (6x+10) = 2x

OpenStudy (anonymous):

sorry *solve the square root

OpenStudy (raden):

square to both sides, gives 6x+10=4x^2 or 4x^2-6x-10=0 or 2x^2-3x-5=0 (2x-5)(x+1)=0 for zeroes, 2x-5=0 --->x1=5/2 x+1=0 --->x2=-1 but, recheck again which the value of x, satisfied to the original problem sqrt(6x+10)=2x for, x=5/2 gives sqrt(6*5/2 + 10) = 2*5/2 sqrt(15+10) = 10/2 sqrt(25) = 5 5 = 5 because LHS equal by RHS, so it is correct as the answer now, check for x=-1 sqrt(6x+10)=2x (just plug x=-1) sqrt(6*(-1) + 10) = 2*(-1) sqrt(-6+10) = -2 sqrt(4) = -2 2 = -2 because LHS not equal by RHS, so incorrect therefore, the answer just x=5/2 only

OpenStudy (raden):

@EmyleighTeryl , are u there

OpenStudy (anonymous):

yes sorry...thank you

OpenStudy (raden):

welcome